A manufacturing company receives orders for engines from two assembly plants. Plant I needs at least 45 ​engines, and plant II needs at least 32 engines. The company can send at most 140 engines to these assembly plants. It costs $30 per engine to ship to plant I and ​$40 per engine to ship to plant II. Plant I gives the manufacturing company ​$20 in rebates toward its products for each engine they​buy, while plant II gives similar ​$15 rebates. The manufacturer estimates that they need at least ​$1500 in rebates to cover products they plan to buy from the two plants. How many engines should be shipped to each plant to minimize shipping​costs? What is the minimum​ cost?

a.) How many engines should be shipped to each plant to minimize shipping​ costs, subject to the given​ constraints?

The number of engines to send to plant I is ____.

The number of engines to send to plant II is ____.

b.) What is the minimum shipping cost, subject to the given constraints?
Answer #1

let number of engunes ehupped to plant plant Cost to shup engute to plant plant 30 one 40 2 then total Cost 40y We need to sh20x15y1500 100 the shaded region is the feasible region (45, 95) -50 (45, 40) (108, 32) y= 32 (51, 32) x+y= 140 -50 50 100 15Comenpounla: (45,45), (45,40), (5, 32) C108,32) - 40 Nentex C Box+ C4S,) 5150 SI50 (51,32) 3000 5280 The mwimum Cost 3000 whe
Answer #2

what is the minimum shipping cost, subject to the given constraints?


answered by: Austin Stephens